OpenStudy (anonymous):

translate from words to algebraic symbols. a room has one wall that is x feet long. T the total length of this wall together with the adjacent wall is 36feet. using this information, write down an algebraic expression for the length of the second wall. im confused, is it just x =36?

OpenStudy (anonymous):

let x be the first wall and y the second wall x + y = 36 y = 36 - x

OpenStudy (anonymous):

since T = 36 y = T - x

OpenStudy (anonymous):

Okay Thanks!